# Hint of Thought prompting: an explainable and zero-shot approach to reasoning tasks with LLMs

### v6

@@ Abstract
As a way of communicating with users and any LLMs like GPT or PaLM2, prompting becomes an increasingly important research topic for better utilization of LLMs. Although simple prompting performs well on single-step questions, it cannot permanently activate the correct knowledge path for multi-step reasoning tasks. The chain of thought (CoT), which often contains zero-shot CoT and few-shot CoT, is a recently developed prompting method that can explain the reasoning process to the LLM and outperforms simple prompting in three challenging reasoning tasks, including arithmetic, symbolic, and commonsense reasoning. In this paper, we propose a novel hint of thought (HoT) prompting with explainability and zero-shot generalization. First, it is decomposed into the following three steps: explainable sub-questions, logical reasoning, and answer extraction. Second, such three steps are sequentially ordered in the format of step-by-step hints, which can be easily adjusted and explained to different tasks. Finally, experimental results demonstrate that our HoT prompting has a significant advantage on the zero-shot reasoning task compared to existing zero-shot CoT. We did zero-shot experiments on math tasks like GSM8K, ADDSUB, AQUA, SVAMP and commonsense tasks such as StrategyQA. In particular, the accuracy of the proposed HoT prompting is improved with GSM8K from 40.50% to 67.80%, with AQUA from 31.9% to 46.4%, with SVAMP from 63.7% to 76.9%, and with ADDSUB from 74.7% to 87.34%, respectively, which even defeats the competitive PoT approach on GSM8k, AQUA, and SVAMP.
@@ 1 Introduction
Many researchers indicate that scaling up the size of generative language models and training datasets plays a critical role in recent NLP research. There are some powerful examples of large language models (LLMs), such as ChatGPT leiter2023chatGPT, PaLM Chowdhery et al. (2022), and LAMDA Thoppilan et al. (2022). Indicated by the robustness of GPT-4 OpenAI (2023), many valuable applications can be developed, and Bubeck et al. (2023) believes that it has the potential to lead the world of artificial intelligence (AI) to the world of artificial general intelligence (AGI).
@@
Although there is a fantastic performance that LLMs can solve single-step or intuitive tasks very well with task-specific zero-shot prompting or in-context learning prompting Liu et al. (2021), models with scale up to 100B+ parameters are struggling to solve multi-step reasoning tasks Rae et al. (2022). To address this drawback, researchers have introduced a method called chain of thought (CoT) prompting Wei et al. (2022a), rather than standard question and answer examples prompting, and it feeds LLMs with a chain of reasoning examples (example). In other words, CoT demonstrates a reasoning path that composes an original complex question into multiple more straightforward steps. With CoT, the performance of language models on a large scale has dramatically improved. For example, a giant PaLM with 540B parameter Chowdhery et al. (2022) with CoT can increase the accuracy of math calculations.
A few-shot prompting with CoT Wei et al. (2022a) or without CoT Brown et al. (2020) can finish many task-specific prompting tasks. Decomposed Prompting Khot et al. (2023) proposed a model to decompose the task into sub-tasks and solve them iterative; however, it requires few-shot prompts and has to prompt a few times to get the result. To do zero-shot reasoning, researchers proposed zero-shot CoT, showing that LLMs are good zero-shot reasoners by simply adding "let’s think step by step" Kojima et al. (2022). However, the solving process is provided by the LLM without any explanation between each step. Also, Program of Thought (PoT) Chen et al. (2023) performed better using Python as an extended tool for zero-shot math reasoning but mainly focused on math reasoning tasks. Both works show the difficulty of reasoning tasks with LLMs. To build an explainable, logical, and user-friendly end-to-end zero-shot prompt method, we propose hint of thought (HoT) from another point of view. To approach the solution, we take the traditional zero-shot CoT as an activation path that activates the correct reasoning path within LLMs. We propose a hints chain that is explainable to humans and efficient for the LLM to understand. It contains three main steps (1) ask for five subquestions; (2) ask pseudocode as an answer for subquestions; (3) get the answer in a wanted way. The hints chain automatically leads the LLM to do zero-shot reasoning by asking small sub-questions and answering sub-questions in pseudocode. The sub-questions can help users understand the “mind map” of the LLM, and pseudocodes provide a transparent, logical reasoning process. Moreover, answering in pseudocode can avoid semantic ambiguity. We used GPT-3.5-turbo Brown et al. (2020) as our LLM; it is designed to give a complete sentence as an answer Ye et al. (2023).
@@ 2.1 Large language models and prompting
A language model (LM) is a model that is designed for estimating the probability distribution of text. In recent research, they found that scaling up the model size can help improve the performance (from a few million Merity et al. (2016) to hundreds of millions Devlin et al. (2019) to hundreds of billions Brown et al. (2020) parameters). And the training data also becomes more extensive, e.g., webtest corpora Gao et al. (2020). These improve the abilities of pre-trained LLMs in many downstream NLP tasks. Unlike the classic paradigm of "pre-train and fine-tune," an LLM that scales to 100B+ parameters displays the ability to few-shot learning Brown et al. (2020). Using in-context learning, we can use prompts to strongly lead the generation to output a desired answer to a specific task. It is called pre-train and prompts Liu et al. (2021).
@@
Based on the drawback of few-shot CoT that costs time and people to design the prompt, Kojima et al. (2022) proposed a zero-shot CoT prompt. They added "Let’s think step by step" or a similar text. This work showed that the potential of LLMs is zero-shot. Compared with standard zero-shot and zero-shot CoT, the latter significantly improved with GSM8K from 12.5% to 40.5%.
2.3 Program of Thought Chen et al. (2023)
To make more accurate calculations on math tasks. PoT proposed the use of an extended tool to calculate the answer. In their work, they ask the LLM to propose and execute a Python code to get the math answer. In their zero-shot experiment, PoT performs better than zero-shot CoT on GSM8K (57.0%), AQuA (43.9%), SVAMP (70.8%), TabMWP (66.5%), MultiArith (92.2%), and Avg (66.1%).
@@ 3 HoT
We propose an explainable, logical, and end-to-end hint of thought prompting (HoT) method that uses reasoning paths to represent knowledge to activate more profound reasoning within an LLM. In our work, an end-to-end hints chain (in the step-by-step format) is used as the input of LLMs, and the LLMs itself is required to generate explainable semantic-level step-by-step reasoning steps without any extra prompts. This employs a chain-like instruction to activate more profound knowledge within the LLM. Consequently, the input prompting to the LLM is "$[X]$. Hints: $hints\_chain$" where $[X]$ is the question, and $hints\_chain$ is our HoT prompting.
3.1 End-to-end hints chain
The hints chain is a chain of instructions that delivers the LLM with instructions step-by-step. The purpose of the hints chain is to tell the LLM how to generate a reasoning chain. There are three parts of such hints chains: (1) Explainable sub-questions: ask the LLM to partition the question into five step-by-step sub-questions; (2) Logical reasoning: tell the LLM to use the step-by-step sub-questions to generate a pseudocode and calculate the answer; (3) Answer extraction: ask the LLM to give a customized format of the answer. The result produced by the hints chain follows the order of these three parts and provides a consequential reasoning chain that supports its answer. It is an end-to-end way, meaning we only need to prompt once for a single question to get an answer, which can help the LLM respond quickly.
@@ 3.2 Explainable question decomposition and sub-question-wise problem solving
Although the prompt in LLM is also in the back box representation, our HoT can produce an understandable and explainable consequent problem-solving procedure. It indeed helps the users to verify the answer more intuitively. Sub-questions not only help the LLM to solve the whole problem but also transparently explain the mind map of the LLM. In our examples (Appendix), the reasoning process and logical reasoning can be clearly seen.
@@ 3.3 Logical pseudo-code reasoning
An interpretable problem-solving chain can enhance LLM to solve reasoning problems more accurately. However, it is likely to have semantic ambiguity in semantic-level reasoning, although LLM is designed to work with language. Therefore, pseudocode, regarded as a more accurate logical language in programming design, is further exploited.
@@ 4.1 Tasks and datasets
We evaluate the HoT prompting on the five datasets on the four main arithmetic reasoning tasks: GSM8K, AQUA, SVAMP, and ADDSUB. In addition, we complete experiments on the big commonsense reasoning benchmark StrategyQA. Note that all the datasets we utilize in our experiments are publicly released.
Dataset # of samples
@@ 4.2 Arithmetic tasks
For the arithmetic reasoning task, the following four benchmarks are considered: i.e., 1) GSM8K Cobbe et al. (2021), 2) AQUA-RAT Ling et al. (2017), 3) SVAMP Patel et al. (2021), and 4) AddSub Hosseini et al. (2014). The first three datasets, especially GSM8K, are more recently published benchmarks and have more challenges because they require multi-step reasoning to solve problems. Also, the AQUA is a multiple-choice question dataset. The baseline results are quoted from zero-shot CoT Kojima et al. (2022) and zero-shot experiment done in PoT Chen et al. (2023). We can observe that HoT has the best performance among the zero-shot reasoning prompts. We design the HoT prompts here, requiring a numeral answer in the third part of the hints chain. From the results, it is easy to see that HoT has a higher potential for arithmetic reasoning than the previous zero-shot approaches.
@@
HoT (Ours) 67.80% 46.4%
Table 2: Results on GSM8K and AQUA
@@ 4.3 Commonsense tasks
We employ StrategyQA Geva et al. (2021) for the commonsense reasoning task, a large dataset that requires the model to conduct implicit multi-hop reasoning to answer questions. The baseline results are from zero-shot CoT.Kojima et al. (2022). The reasoning accuracy with HoT is about 30% higher than CoT. We make the HoT prompt here, which requires a Yes/No in the third part of the hints chain. It is readily observed from the results that HoT has a higher potential for commonsense reasoning than existing zero-shot approaches.
@@ 4.4 Answer verification
GPT-3.5-turbo always answers (output) in a complete sentence, making it difficult to automatically verify if the output is the same answer as that in the metrics. To check the answer, we independently prompt again to the LLM with its answer as a record. In this verification prompt, we take the model answer as user input and verify if it’s the same as the answer given by the LLM.
Figure 2: Full pipeline of answer verification
To verify the reliability of our verification method, we produce a correct validation and a wrong validation by manually setting up the positive and negative samples in GSM8k which is the largest dataset within our experiment. We manually set a prompt that always with the fake/correct answer: " Therefore, the answer is fake/correct_answer" and then we follow our verification process and gain a result. In the correct validation test, we obtain 1293 true positive samples out of 1319 samples. In the wrong validation test, we obtain 1299 false negative samples out of 1319 samples. So, the correct validation rate is 98.02%, and the wrong validation rate is 98.48%. This shows the reliability of our verification method is strong and robust.
@@ 5 Error Analysis
We divide the reasoning errors into (1) reasoning errors and (2) calculation errors. Our experimental results illustrate that reasoning errors always occur when the question (Error on GSM8K) has semantic ambiguity or if it is too complex. It is hard to avoid such errors because it is hard even for a human being. Although the calculation error rarely occurs, this can happen, and this can be potentially tackled by adding an external calculator.
@@ 6 Ablation study
6.1 The number of sub-questions.
We conduct HoT experiments with three sub-questions (HoT-3) and seven sub-questions (HoT-7) on the SVAMP dataset, respectively, while the standard one is HoT-5 with five sub-questions. We can see that more sub-questions may not provide a more accurate answer. The LLMs start making stories when those steps are not needed. Also, more sub-questions cost more for the LLMs to respond, which takes more time and effort.
Method SVAMP
zero-shot CoT 52.3%
PoT with zero-shot 43.9.0%
HoT-3 46.1%
HoT-5 46.4%
HoT-7 44.4%
Table 5: Ablation study on SVAMP
@@ 7.1 Complex reasoning with LLMs
Reasoning skills are essential for general intelligence systems, and the ability to reason in LLMs gained significant attention from the research community. Several studies Brown et al. (2020); Smith et al. (2022) have shown that asking pre-trained models to produce step-by-step reasoning or fine-tuning Cobbe et al. (2021) can increase their ability on complex reasoning tasks. GPT-3 Brown et al. (2020) has illustrated its robust few-show reasoning Wei et al. (2022b); Wang (2022); Chowdhery et al. (2022) that a few examples in natural language are given to the model to describe the task. The most classic reasoning tasks are mathematical reasoning. PoT Chen et al. (2023) has shown great ability on math reasoning tasks with LLMs with the help of Python programs. They aim to generate an executive Python program by the LLM to solve math problems. However, their work primarily focuses on math reasoning tasks. A more general approach would be CoT Wei et al. (2022a), which works well on mathematical, logical, common sense, and symbolic reasoning tasks with few-shot prompts.
@@ 7.2 Zero-shot reasoning with LLMs
It was indicated that LLMs have excellent zero-shot abilities in many system-1 tasks, including reading comprehension, translation, and summarization Radford et al. (2019). This ability can also be fine-tuned to get a better performance Ouyang et al. (2022). However, we focus on system-2 tasks beyond system-1 tasks. The recent work, zero-shot CoT, increases zero-shot performance. Also, PoT in zero-shot format provides good results in math reasoning tasks.

### v7

@@ Abstract
Prompting becomes an increasingly important research topic for better utilization of LLMs. Although simple prompting performs well on single-step questions, it cannot permanently activate the correct knowledge path for multi-step reasoning tasks. The chain of thought (CoT), which often contains zero-shot CoT and few-shot CoT, is a recently developed prompting method that can explain the reasoning process to the LLM and outperforms simple prompting in three challenging reasoning tasks, including arithmetic, symbolic, and commonsense reasoning. Inspired by zero-shot CoT, and further extending the zero-shot ability, this paper proposes a novel hint of thought (HoT) prompting with explain-ability and zero-shot logicalbility. It is decomposed into three steps: explainable sub-questions, logical reasoning, and answering. Such three steps are sequentially ordered in step-by-step hints, which can be easily adjusted and explained to different tasks. Finally, experimental results demonstrate that our HoT prompting has a significant advantage on the zero-shot reasoning task compared to zero-shot CoT. We did zero-shot experiments on math tasks like GSM8K, ADDSUB, AQUA, SVAMP, and commonsense tasks such as StrategyQA. In particular, the accuracy of the proposed HoT prompting is improved with GSM8K from 40.50% to 70.65%, with AQUA from 31.9% to 46.4%, with SVAMP from 63.7% to 76.9%, and with ADDSUB from 74.7% to 87.34%, respectively, which even defeats the competitive PoT approach on GSM8k, AQUA, and SVAMP.
@@ 1 Introduction
Many researchers indicate that scaling up the size of generative language models and training datasets plays a critical role in recent NLP research. There are some powerful examples of large language models (LLMs), such as ChatGPT Leiter et al. (2023), PaLM Chowdhery et al. (2022), and LAMDA Thoppilan et al. (2022). Indicated by the robustness of GPT-4 OpenAI (2023), many valuable applications can be developed, and Bubeck et al. (2023) believes that it has the potential to lead the world of artificial intelligence (AI) to the world of artificial general intelligence (AGI).
@@
Although there is a fantastic performance that LLMs can solve single-step or intuitive tasks very well with task-specific zero-shot prompting or in-context learning prompting Liu et al. (2021), models with scale up to 100B+ parameters are struggling to solve multi-step reasoning tasks Rae et al. (2022). To address this drawback, researchers have introduced a method called chain of thought (CoT) prompting Wei et al. (2022), rather than standard question and answer examples prompting, and it feeds LLMs with a chain of reasoning examples. In other words, CoT demonstrates a reasoning path that composes an original complex question into multiple more straightforward steps. With CoT, the performance of language models on a large scale has dramatically improved. For example, a giant PaLM with 540B parameter Chowdhery et al. (2022) with CoT can increase the accuracy of math calculations.
A few-shot prompting with CoT Wei et al. (2022) or without CoT Brown et al. (2020) can finish many task-specific prompting tasks. Decomposed Prompting Khot et al. (2023) proposed a model to decompose the task into sub-tasks and solve them iterative; however, it requires few-shot prompts and has to prompt a few times to get the result. To do zero-shot reasoning, researchers proposed zero-shot CoT, showing that LLMs are good zero-shot reasoners by simply adding "let’s think step by step" Kojima et al. (2022). However, the solving process is provided by the LLM without any explanation between each step. Also, Program of Thought (PoT) Chen et al. (2023) performed better using Python as an extended tool for zero-shot math reasoning but mainly focused on math reasoning tasks. Both works show the difficulty of zero-shot reasoning tasks with LLMs. To build an explainable, logical, and flexible zero-shot prompt method, we propose the hint of thought (HoT) prompting. To approach the solution, we take the traditional zero-shot CoT as an activation path that activates the correct reasoning path within LLMs. We propose a hint chain that is explainable to humans and efficient for the LLM to understand. It contains three main parts (1) ask for five subquestions; (2) ask pseudocode as an answer for subquestions; (3) get the answer in a wanted way. The hint chain automatically leads the LLM to do zero-shot reasoning by asking small sub-questions and answering sub-questions in pseudocode. The output of the sub-questions can help users understand the “mind map” of the LLM, and pseudocodes provide a more transparent and logical reasoning process.
Moreover, answering in pseudocode can avoid semantic ambiguity. Because of this advantage, HoT can give a more precise reasoning process than zero-shot CoT. We used GPT-3.5-turbo Brown et al. (2020) as our LLM, a fine-tuned question-answering model. We chose this model because it is a chat model, and it is more likely to be used as a chatbot, which is relatively weaker in specific reasoning tasks. However, HoT can ideally use its conversation ability to interpret the question and increase its reasoning ability to avoid its weakness.
@@ 2.1 Large language models and prompting
A language model (LM) is a model that is designed to estimate the probability distribution of text. In recent research, they found that scaling up the model size can help improve the performance (from a few million Merity et al. (2016) to hundreds of millions Devlin et al. (2019) to hundreds of billions Brown et al. (2020) parameters). And the training data also becomes more extensive, e.g., webtest corpora Gao et al. (2020). These improve the abilities of pre-trained LLMs in many downstream NLP tasks. Unlike the classic paradigm of "pre-train and fine-tune," an LLM that scales to 100B+ parameters displays the ability to few-shot learning Brown et al. (2020). Using in-context learning, we can use prompts to strongly lead the generation to output a desired answer to a specific task. It is called pre-train and prompts Liu et al. (2021).
@@
Based on the drawback of the few-shot CoT that costs time and people to design the prompt, Kojima et al. (2022) proposed a zero-shot CoT prompt. They added "Let’s think step by step" or a similar text. This work showed that the potential of LLMs is zero-shot. Compared with standard zero-shot and zero-shot CoT, the latter significantly improved with GSM8K from 12.5% to 40.5%. However, "Let’s think step by step" is a settled prompt that can not be flexible in every reasoning task. And this inspired us to invent a more flexible zero-shot prompt.
2.3 Program of Thought (PoT) Chen et al. (2023)
To make more accurate calculations on math tasks. PoT proposed the use of an extended tool to calculate the answer. In their work, they ask the LLM to propose and execute a Python code to get the math answer. In their zero-shot experiment, PoT performs better than zero-shot CoT on GSM8K (57.0%), AQuA (43.9%), SVAMP (70.8%), TabMWP (66.5%), MultiArith (92.2%), and Avg (66.1%). However, this program-based work does not interpret the answers well. And this inspired us to invent a more interpretable prompt.
@@ 3 HoT
To increase the flexibility of zero-shot CoT and the interpretation of PoT. We propose an explainable and logical hint of thought (HoT) prompting that uses reasoning paths to represent knowledge to activate more profound reasoning within an LLM. In our work, a hint chain (in the step-by-step format) is used as the input of LLMs, and the LLMs themselves are required to generate explainable semantic-level step-by-step reasoning steps without any extra prompts. This employs a chain-like instruction to activate more profound knowledge within the LLM. Consequently, the input prompting to the LLM is "$[X]$. Hints: $hint\_chain$" where $[X]$ is the question, and $hint\_chain$ is our HoT prompting.
3.1 Flexible Hint chain
The hint chain is a chain of instructions that delivers the LLM with instructions step-by-step. The purpose of the hint chain is to tell the LLM how to generate a reasoning chain. There are three parts of such hint chains: (1) Explainable sub-questions: ask the LLM to partition the question into five step-by-step sub-questions; (2) Logical reasoning: tell the LLM to use the step-by-step sub-questions to generate a pseudocode and calculate the answer; (3) Answering: ask the LLM to give a customized format of the answer. The result produced by the hint chain follows the order of these three parts and provides a consequential reasoning chain that supports its answer.
@@ 3.2 Explainable question decomposition and sub-question-wise problem solving
Although the LLM is in the back box representation, our HoT can produce an understandable and explainable consequent problem-solving procedure. It indeed helps the users verify the answer more intuitively. Sub-questions help the LLM solve the whole problem and transparently explain the LLM’s mind map. In our work, a intuitive and logical reasoning process can be seen.
@@ 3.3 Logical pseudo-code reasoning
An interpretable problem-solving chain can enhance LLM’s ability to solve reasoning problems more accurately. However, although LLM is designed to work with language, it is likely to have semantic ambiguity in semantic-level reasoning. Therefore, pseudocode, regarded as a more accurate logical language in programming design, is used in HoT and this produce a higher reasoning accuracy.
@@ 4.1 Tasks and datasets
We evaluate the HoT prompting on five datasets for the four main arithmetic reasoning tasks: GSM8K, AQUA, SVAMP, and ADDSUB. In addition, we complete experiments on the big commonsense reasoning benchmark StrategyQA. Note that all the datasets we utilize in our experiments are publicly released. And LLMs have a randomness property, and the extract accuracy might not always be reproduced.
Table 1: Number of samples in different datasets
Dataset Number of samples
@@ 4.2 Arithmetic tasks
For the arithmetic reasoning task, the following four benchmarks are considered: i.e., 1) GSM8K Cobbe et al. (2021), 2) AQUA-RAT Ling et al. (2017), 3) SVAMP Patel et al. (2021), and 4) AddSub Hosseini et al. (2014). The first three datasets, especially GSM8K, are more recently published benchmarks and have more challenges because they require multi-step reasoning to solve problems. Also, the AQUA is a multiple-choice question dataset. The baseline results are quoted from zero-shot CoT Kojima et al. (2022) and zero-shot experiment done in PoT Chen et al. (2023). We can observe that HoT performs the best among the zero-shot reasoning prompts with GPT3s. We design the HoT prompts here, requiring a numeral answer in the third part of the hint chain. From the experiment results, it is easy to see that HoT has a higher potential for arithmetic reasoning than the previous zero-shot approaches.
Table 2: Results on GSM8K and AQUA
@@
HoT (Ours) 70.65% 46.4%
Table 3: Results on SVAMP and ADDSUB
@@ 4.3 Commonsense tasks
We employ StrategyQA Geva et al. (2021) for the commonsense reasoning task, a large dataset that requires the model to conduct implicit multi-hop reasoning to answer questions. The baseline results are from zero-shot CoT Kojima et al. (2022). The reasoning accuracy with HoT is about 30% higher than CoT. We make the HoT prompt here, which requires a Yes/No in the third part of the hint chain. It is readily observed from the results that HoT has a higher potential for commonsense reasoning than the existing zero-shot prompt approach.
Table 4: Results on StrategyQA
@@ 4.4 Answer verification
GPT-3.5-turbo always answers (output) in a complete sentence, making it difficult to automatically verify if the output is the same answer as that in the metrics. To check the answer, we independently prompt the LLM again with its answer as a record. In this verification prompt, we take the model answer as user input and verify if it’s the same as the answer given by the LLM.
To verify the reliability of our verification method, we produced a correct and wrong validation by manually setting up the positive and negative samples in GSM8k, the largest dataset within our experiment. We manually set a prompt that always has the fake or correct answer: " Therefore, the answer is fake or correct_answer" and then we follow our verification process and gain a result. In the correct validation test, we obtain 1293 true positive samples out of 1319 samples. We obtain 1299 false negative samples in the wrong validation test out of 1319. So, the correct validation rate is 98.02%, and the wrong validation rate is 98.48%. This shows the reliability of our verification method is strong.
@@ 5 Error Analysis
We divide the reasoning errors into (1) reasoning errors and (2) calculation errors. Our experimental results illustrate that reasoning errors always occur when the question has semantic ambiguity or if it is too complex. Avoiding such errors is difficult because it’s challenging even for a human.
@@ 6 Ablation study
We remove the subquestions and pseudocode parts separately and perform a GSM8K test. The result shows that HoT with only subquestions performs similarly to zero-shot CoT in terms of accuracy and interoperability. Meanwhile, HoT with only pseudo-code produces accuracy similar to that of HoT but with less interoperability with only outputting pseucdocodes. This experiment shows that the pseudocode part gives the logical reasoning ability of HoT, and the subquestions part gives interoperability.
Table 5: Ablation study on GSM8K
Method Accuracy
HoT 70.65%
HoT without pseudo-code 42.3%
HoT without subquestions 69.2%
@@ 7.1 Complex reasoning with LLMs
Reasoning skills are essential for general intelligence systems, and the ability to reason in LLMs gained significant attention from the research community. Several studies Brown et al. (2020); Cobbe et al. (2021) have shown that asking pre-trained models to produce step-by-step reasoning or fine-tuning can increase their ability on complex reasoning tasks. GPT-3 Brown et al. (2020) has illustrated its robust few-shot reasoning that a few examples in natural language are given to the model to describe the task. The most classic reasoning task is mathematical reasoning. PoT Chen et al. (2023) has shown great ability on math reasoning tasks with LLMs with the help of Python programs. They aim to generate an executive Python program by the LLM to solve math problems. However, their work primarily focuses on math reasoning tasks. A more general approach would be CoT Wei et al. (2022), which works well on mathematical, logical, common sense, and symbolic reasoning tasks with few-shot prompts.
@@ 7.2 Zero-shot reasoning with LLMs
It was indicated that LLMs have excellent zero-shot abilities in many system-1 tasks, including reading comprehension, translation, and summarization Radford et al. (2019). This ability can also be fine-tuned to get a better performance. However, we focus on system-2 tasks beyond system-1 tasks. The recent work, zero-shot CoT, increases zero-shot performance. Also, PoT in zero-shot format provides good results in math reasoning tasks.